OpenStudy (anonymous):

a ladder 15 feet long leans against the wall and forms an angle of 45 degrees with the ground. how far from the wall is the ladder?

OpenStudy (whpalmer4):

 use Pythagorean theorem...

OpenStudy (whpalmer4):

smart-alec answer: "it's touching!" :-)

OpenStudy (austinl):

darn it, you beat me to the drawing.

OpenStudy (whpalmer4):

presumably the problem really wants to know the distance the base of the ladder is from the wall...

OpenStudy (austinl):

You could also use SohCahToa as well, I actually prefer that method. but your way should give an exact answer.

OpenStudy (anonymous):

i did the therom and i cant get the answer still

OpenStudy (austinl):

\[h^2+h^2=c^2\] From that you should be able to solve for h, which is the distance from the ladder to the wall.

OpenStudy (whpalmer4):

Pythagorean theorem says the sum of the squares of the sides = the square of the hypotenuse. Here we have the hypotenuse = 15, and the two sides are equal (because the 45 degree angle implies an isosceles triangle). \[h^2 + h^2 = 15^2\]\[2h^2 = 225\]\[h = \sqrt{225}/\sqrt{2} = \]

OpenStudy (whpalmer4):

it is worth remembering that the ratio of the sides in a right isosceles triangle is \(1:1:\sqrt{2}\)

OpenStudy (whpalmer4):

so one could just write down the answer immediately as \(15/\sqrt{2}\) or \(15\sqrt{2}/2\)

OpenStudy (austinl):

Very nice work sir! Kudos.

OpenStudy (whpalmer4):

@minxgirl did you figure out where you went amiss?